OpenStudy (anonymous):

What is the value of x in the proportion (x-1)/5=(4x+2)/35 ? a.1/3 b.1/5 c.3 d.-5/3

hero (hero):

Cross Multiply: 35(x-1) = 5(4x+2) Distribute: 35x - 35 = 20x + 10 Combine Like Terms: 35x - 20x = 10 + 35 Simplify: 15x = 45 Divide to Isolate x: x = 45/15 x = 3